# MeteoRA: Multiple-tasks Embedded LoRA for Large Language Models

### v2

@@ Abstract
The pretrain+fine-tune paradigm is foundational for deploying large language models (LLMs) across various downstream applications. Within this framework, Low-Rank Adaptation (LoRA) stands out for its parameter-efficient fine-tuning (PEFT), producing numerous off-the-shelf task-specific LoRA adapters. However, this approach requires explicit task intention selection, posing challenges for autonomous task sensing and switching during inference with multiple LoRA adapters embedded in a single LLM. In this work, we introduce MeteoRA (Multiple-Tasks embedded LoRA), a scalable and efficient framework that embeds multiple task-specific LoRA adapters into the base LLM via a full-mode Mixture-of-Experts (MoE) architecture. This framework also includes novel MoE forward acceleration strategies to address the efficiency challenges of traditional MoE implementations. Our evaluations, using the LlaMA2-13B and LlaMA3-8B base models equipped with 28 off-the-shelf LoRA adapters through MeteoRA, demonstrate equivalent performance with the traditional PEFT method. Moreover, the LLM equipped with MeteoRA achieves superior performance in handling composite tasks, effectively solving ten sequential problems in a single inference pass, thereby demonstrating the framework’s enhanced capability for timely adapter switching and multi-LoRA fusion.
@@ 1 Introduction
Despite the success of LoRA in the pretrain-fine-tune paradigm, several challenges remain. A primary challenge is the ability of multi-LoRA embedded LLMs to autonomously and on-demand LoRA selection during inference, a process that should allow the LLM to handle different tasks by activating the appropriate LoRA adapters without explicit user instructions. Furthermore, managing composite tasks that require timely switching between LoRA adapters presents difficulties, especially when these tasks involve multiple sub-problems each requiring specific adapter activation. Current approaches such as Huggingface PEFT and S-LoRA, while capable of serving multiple adapters simultaneously, mainly focus on loading rather than autonomously activating adapters based on sensed intentions, thus requiring manual intervention. Similarly, existing LoRA fusion methods such as LoRAHub (Huang et al., 2023) and MoA (Feng et al., 2024), although they integrate and merge knowledge from various adapters, are not suitably designed to fuse a wide range of existing LoRA adapters with such a limited MoE framework, and lack the evidence in effectively managing dynamic adapter switching during inference for composite tasks.
@@
In this paper, we introduce a novel multi-tasks embedded LoRA framework for LLMs to solve these challenges. The framework proposes a MoE-style module called MeteoRA. Each MeteoRA module provides a trainable Gating network with MoE forward acceleration strategies (overcome the efficiency issue in naive MoE, especially when number of experts is much larger than 8) for all LoRA adapters’ low-rank matrices in the linear layer. As shown in Figure 1, the MeteoRA module is applicable for all kinds of layers in Transformer-based LLMs (Q, K, V, and O in attention module and up_proj, gating for SiLU (Elfwing et al., 2018), and down_proj in MLP Module). Through fine-tuning all gates with minimal resources, MeteoRA effectively integrates all embedded LoRA adapters into the base LLM model with the ability of autonomously on-demand LoRA selection, without the requirements of any explicit user or system instructions. Furthermore, the presence of numerous gates111For the LlaMA3-8B model, there are 224 MeteoRA modules in total, with each of the 32 decoder layers containing 7 gates (Q, K, V, O, up_proj, gating, and down_proj). enhances the model with a full-mode MoE architecture, showing the capability of timely LoRA switching, addressing composite tasks with only two-shot examples as illustrations for all inputs. Our empirical evaluations, which embedded 28 LoRA adapters with MeteoRA to LlaMA2-13B-base and LlaMA3-8b-base, highlight the full-mode MoE capabilities and demonstrate a significant performance enhancement. This improvement is particularly notable in handling composite tasks, showcasing the efficacy of the MeteoRA framework.
The primary contributions of MeteoRA are summarized as follows:
@@
Scalable LoRA integration: MeteoRA framework for multi-LoRA embedding advances the LLM’s capability of autonomous on-demand LoRA selection and timely LoRA switching.
@@
Organization. Section 2 discusses the background of LoRA adapters and the mixture-of-experts. Section 3 formulates and details the proposed method for embedding massive LoRA adapters into one base LLM. Section 4 presents empirical evaluations. Section 5 covers related work. Section 6 discusses the limitations. Section 7 concludes the paper.
@@ 3.1 MeteoRA architecture
Given a base LLM and LoRA adapters that are fine-tuned with the distinct tasks on separately, our objective is to integrate LoRA adapters into the base via MeteoRA framework, resulting in a LoRA embedded model . Figure 1 demonstrates the MeteoRA module complemented to each basic linear layer in LLM. Assume that the given LoRA adapters are off-the-shelf components, available from open-source communities or pre-fine-tuned for specific tasks, and each LoRA adapter contains a set of low-rank matrices . MeteoRA furnishes each basic linear layer with a wide MoE architecture to embed the low-rank matrices provided by LoRA adapters.
Figure 2 shows the architecture of MeteoRA module. To embed LoRA adapters, MeteoRA module leverages the MoE architecture by injecting a trainable Gating network together with pairs of to . By applying , MeteoRA selects pairs of with the top- highest gated weights for each . It then proceeds with the forward pass as follows:
@@
where denotes the unnormalized gated logits for the -th LoRAs. By doing this, the Gating network performs as a routing strategy for selecting the appropriate LoRA adapters based on the layer’s input. Each MeteoRA module contains a Gating network, and the Gating networks from different MeteoRA modules make decisions based on their own inputs, the selection of LoRA adapters could be dynamically switched and adjusted in the forward process of each MeteoRA module through all LLM’s decoder blocks. MeteoRA also applies top- and top- gating strategies as detailed in Appendix A.1.
@@ 4 Evaluation
We conduct experiments on individual and composite tasks as detailed in Section 4.1. For our base models, we use two well-known LLMs, LlaMA2-13B (Touvron et al., 2023) and LlaMA3-8B (Meta, 2024). The code and the models are available222The implementation code is accessible at https://github.com/ParagonLight/meteor-of-lora.
@@ 4.1 Evaluation settings
Metrics. We apply a zero-shot evaluation setting for all tasks, adding brief task descriptions for tasks such as CNN/DailyMail and the three translation tasks that do not inherently include task descriptions. As for metrics, we use accuracy for multiple-choice tasks and GSM8K while employing metrics such as BLEU, ROUGE-1, ROUGE-1, and ROUGE-L for other tasks.
Models. We use LlaMA2-13B and LlaMA3-8B as the base LLMs for LoRA and MeteoRA adaption. Both LlaMA models are pretrained LLMs and do not include the process of instruction tuning. We train specific LoRA adapters for each task using their respective training sets. The process of training LoRA adapters could be offline or dismissed when off-the-shelf LoRA is accessible. Then, the Gating networks, which embed the adapters in the MeteoRA module, are fine-tuned efficiently based on the balanced dataset. The Gating networks for 28 tasks takes no more than 10 hours to reach the convergence with 4 H800 training via Accelerate (Gugger et al., 2022).
For baseline comparisons, we train one LoRA adapter (i.e., LoRA-F) using a mixed training set from all 28 tasks, and another LoRA adapter (i.e., LoRA-B) with the balanced dataset designed for training the Gating network. We also use Huggingface PEFT (short in PEFT) loading all 28 LoRA adapters (same ones used for MeteoRA) with explicit LoRA activation information during evaluation as a reference model. For details, all LoRA adapters interact with all seven linear layers in LLaMA’s Decoder layer, configured with , , and a learning rate of . Due to some tasks having small training sets, the batch size for fine-tuning is set to . All our experiments were conducted on a GPU server with five H800 80G GPUs. Notice that we carefully selected the training hyperparameters for the LoRA-F and LoRA-B to ensure that their performance on the 28 tasks would not be excessively incomparable.
@@ 4.2 Main results
Figures 3(a) and 3(b) demonstrate the performance of the MeteoRA models, the compared methods (LoRA-F and LoRA-B), and a reference model PEFT based on LlaMA2-13B and LlaMA3-8B, respectively, across the selected 28 tasks. Table 1 shows the averaged scores in various matrics for all methods.
@@
| MeteoRA (top-1) | 0.748 / 0.806 | 41.74 / 44.98 | 0.336 / 0.337 | 0.160 / 0.158 | 0.314 / 0.313 |
| MeteoRA (top-2) | 0.754 / 0.811 | 36.73 / 45.64 | 0.336 / 0.338 | 0.160 / 0.158 | 0.313 / 0.314 |
| LoRA-F | 0.729 / 0.767 | 41.27 / 42.93 | 0.318 / 0.327 | 0.136 / 0.157 | 0.294 / 0.306 |
| LoRA-B | 0.665 / 0.750 | 37.98 / 38.47 | 0.318 / 0.343 | 0.128 / 0.171 | 0.288 / 0.320 |
The evaluation results indicate that, regardless of the base LLM, the MeteoRA models utilizing the top-1 strategy achieve performance very close to the reference model PEFT, while no explicit LoRA activation/deactivation is required in MeteoRA. Although LLMs with both LoRA-F and LoRA-B reach comparable performance on several certain tasks, they exhibit significantly poorer outcomes on others. Additionally, MeteoRA employing the top-2 strategy, despite occasionally showing greater capability loss compared to MeteoRA with top-1 strategy, occasionally outperforms PEFT with adapters trained directly on the individual tasks. This suggests that the component in the loss function (Equation 4) becomes influential in these cases, indicating a beneficial mix of LoRA adapters from various tasks for future study.
@@ 4.4 Efficiency
To assess the efficiency of our novel forward pass designs using custom GPU kernel operators, we truncate samples from each test dataset of all 28 tasks. We evaluate these designs alongside four variants with the same hyperparameters: the upper-bound single-lora, the baseline loop-original, and two novel forward acceleration strategies based on : bmm-torch and bmm-triton, implemented by PyTorch and Triton respectively. Figure 5 displays the histogram of the overall root-of-runtime metric for each task and design. Additional efficiency evaluation experiments are detailed in Appendix A.7.
@@ 5 Related Work
Multi-task fusion. Our proposed method falls into the field of LoRA adapter composition for multi-task fusion. The first category focuses on fusing the entire models. Researchers mainly study model ensembling and multi-task learning to achieve this goal. Existing works integrate the models under the setting of shared model architecture (Matena and Raffel, 2022; Jin et al., 2022; Wu et al., 2023c; Yadav et al., 2024). Others focus on merging models with various architectures or from different tasks. Both methods (Stoica et al., 2023; Liu et al., 2022) try to merge models that are trained for various tasks without additional training. The second category is more concerned with fusion in terms of the tasks. Ilharco et al. (2022) proposes a model editing method via task vectors. Sun et al. (2022) leverage in-context learning with few-shots to enhance the performance of unseen tasks. However, these methods require multi-task training or prior knowledge for the evaluation tasks. Our method embeds off-the-shelf LoRA adapters with a Gate network in the MeteoRA module. None of the examples (zero-shot) are required for all individual tasks.
@@ A.4 Experimental results of 28 tasks
Table 4, Table 5, Table 6 and Table 7 show the detailed evaluation results of different models on the 28 selected tasks. When drawing Figure 3, for tasks we use BLEU and ROUGE as metrics, we selected BLEU for news_commentary_de, news_commentary_es, and news_commentary_it, while opting for ROUGE-L for the remaining tasks.
@@
| Task Name | PEFT (reference) | MeteoRA (top-1) | MeteoRA (top-2) | LoRA-F | LoRA-B |
| abstract_narrative_understanding | 0.863 | 0.860 | 0.858 | 0.758 | 0.720 |
| contextual_parametric_knowledge_conflicts | 0.999 | 0.999 | 0.999 | 0.999 | 0.994 |
| cs_algorithms | 0.841 | 0.818 | 0.841 | 0.848 | 0.818 |
| disfl_qa | 0.690 | 0.684 | 0.679 | 0.670 | 0.573 |
| elementary_math_qa | 0.801 | 0.725 | 0.794 | 0.671 | 0.375 |
| epistemic_reasoning | 1.000 | 0.998 | 1.000 | 1.000 | 0.995 |
| formal_fallacies_syllogisms_negation | 0.999 | 0.996 | 0.999 | 0.921 | 0.565 |
| goal_step_wikihow | 0.906 | 0.872 | 0.887 | 0.877 | 0.842 |
| gsm8k | 0.458 | 0.439 | 0.420 | 0.428 | 0.338 |
| language_identification | 0.874 | 0.854 | 0.872 | 0.728 | 0.542 |
| logical_deduction | 0.720 | 0.717 | 0.713 | 0.653 | 0.680 |
| object_counting | 0.740 | 0.725 | 0.735 | 0.690 | 0.725 |
| paragraph_segmentation | 0.214 | 0.182 | 0.195 | 0.274 | 0.214 |
| play_dialog_same_or_different | 0.649 | 0.649 | 0.649 | 0.649 | 0.650 |
| question_selection | 0.927 | 0.934 | 0.937 | 0.801 | 0.794 |
| reasoning_about_colored_objects | 0.950 | 0.940 | 0.930 | 0.950 | 0.950 |
| strategyqa | 0.731 | 0.722 | 0.742 | 0.729 | 0.722 |
| tracking_shuffled_objects | 0.188 | 0.192 | 0.173 | 0.181 | 0.188 |
| unit_conversion | 0.755 | 0.735 | 0.727 | 0.779 | 0.707 |
| vitaminc_fact_verification | 0.899 | 0.897 | 0.897 | 0.908 | 0.812 |
| winowhy | 0.802 | 0.767 | 0.797 | 0.797 | 0.767 |
| Average | 0.762 | 0.748 | 0.754 | 0.729 | 0.665 |

### v3

@@ Abstract
The pretrain+fine-tune paradigm is foundational for deploying large language models (LLMs) across various downstream applications. Within this framework, Low-Rank Adaptation (LoRA) stands out for its parameter-efficient fine-tuning (PEFT), producing numerous reusable task-specific LoRA adapters. However, this approach requires explicit task intention selection, posing challenges for autonomous task sensing and switching during inference with multiple existing LoRA adapters embedded in a single LLM. In this work, we introduce MeteoRA (Multiple-tasks embedded LoRA), a scalable and efficient framework that reuses multiple task-specific LoRA adapters into the base LLM via a full-mode Mixture-of-Experts (MoE) architecture. This framework also includes novel MoE forward acceleration strategies to address the efficiency challenges of traditional MoE implementations. Our evaluation, using the LlaMA2-13B and LlaMA3-8B base models equipped with 28 existing LoRA adapters through MeteoRA, demonstrates equivalent performance with the traditional PEFT method. Moreover, the LLM equipped with MeteoRA achieves superior performance in handling composite tasks, effectively solving ten sequential problems in a single inference pass, thereby demonstrating the framework’s enhanced capability for timely adapter switching.
@@ 1 Introduction
Despite the success of LoRA in the pretrain-fine-tune paradigm, several challenges remain. When reusing existing LoRA adapters, a primary challenge is the ability of multi-LoRA embedded LLMs to autonomously and on-demand LoRA selection during inference, a process that should allow LLM to handle different tasks by activating the appropriate LoRA adapters without explicit user instructions. Furthermore, managing composite tasks that require timely switching between LoRA adapters presents difficulties, especially when these tasks involve multiple sub-problems each requiring specific adapter activation. Current approaches such as Huggingface PEFT and S-LoRA, while capable of serving multiple existing adapters simultaneously, mainly focus on loading rather than autonomously activating adapters, thus requiring manual intervention. Similarly, current LoRA fusion methods such as LoRAHub (Huang et al., 2023) and MoA (Feng et al., 2024), although they integrate and merge knowledge from various adapters, are not suitably designed to fuse a wide range of existing LoRA adapters with such a limited MoE framework, and lack the evidence in effectively managing dynamic adapter switching during inference for composite tasks.
@@
In this paper, we introduce a novel multi-tasks embedded LoRA framework for LLMs to reuse existing LoRAs with the ability of autonomous task sensing and switching. The framework proposes a MoE-style module called MeteoRA. Each MeteoRA module provides a trainable Gating network with MoE forward acceleration strategies (overcome the efficiency issue in naive MoE, especially when number of experts is much larger than 8) for all LoRAs’ low-rank matrices in the linear layer. As shown in Figure 1, the MeteoRA module is applicable for all kinds of layers in Transformer-based LLMs (Q, K, V, and O in attention module and up_proj, gating for SiLU (Elfwing et al., 2018), and down_proj in MLP Module). Through fine-tuning all gates with minimal resources, MeteoRA effectively integrates the existing LoRA adapters into the base LLM model with the ability of autonomously on-demand LoRA selection, without the requirements of any explicit user or system instructions. Furthermore, the presence of numerous gates111For the LlaMA3-8B model, there are 224 MeteoRA modules in total, with each of the 32 decoder layers containing 7 gates (Q, K, V, O, up_proj, gating, and down_proj). enhances the model with a full-mode MoE architecture, showing the capability of timely LoRA switching, addressing composite tasks with only two-shot examples as illustrations for all inputs. Our empirical evaluations, which embedded 28 existing LoRA adapters with MeteoRA to LlaMA2-13B-base and LlaMA3-8b-base, highlight the full-mode MoE capabilities and demonstrate a significant performance enhancement. This improvement is particularly notable in handling composite tasks, showcasing the efficacy of the MeteoRA framework. The primary contributions of MeteoRA are summarized as follows:
@@
Scalable LoRA integration: MeteoRA framework for reusing existing LoRA adapters advances the LLM’s capability of autonomous on-demand LoRA selection and switching.
@@
Organization. Section 2 discusses the background of LoRA adapters and the mixture-of-experts. Section 3 formulates and details the proposed method for embedding numerous existing LoRA adapters into one base LLM. Section 4 presents empirical evaluations. Section 5 covers related work. Section 6 discusses the limitations. Section 7 concludes the paper.
@@ 3.1 MeteoRA architecture
Given a base LLM and existing LoRA adapters that have already been fine-tuned with the distinct tasks on separately, our objective is to integrate the existing LoRA adapters into the base via MeteoRA framework, resulting in a LoRA embedded model . Figure 1 demonstrates the MeteoRA module complemented to each basic linear layer in LLM. The reused LoRA adapters are off-the-shelf ones, available from open-source communities or have been fine-tuned for specific tasks, Each LoRA adapter contains a set of low-rank matrices . MeteoRA furnishes each basic linear layer with a wide MoE architecture to embed the low-rank matrices provided by LoRA adapters.
Figure 2 shows the architecture of MeteoRA module. To embed existing LoRA adapters, MeteoRA module leverages the MoE architecture by injecting a trainable Gating network together with existing pairs of to . By applying , MeteoRA selects pairs of with the top- highest gated weights for each . It then proceeds with the forward pass as follows:
@@
where denotes the unnormalized gated logits for the -th LoRAs. By doing this, the Gating network performs as a routing strategy for selecting the appropriate LoRA adapters based on the layer’s input. Each MeteoRA module contains a Gating network, and the Gating networks from different MeteoRA modules make decisions based on their own inputs, the selection of LoRA adapters could be dynamically switched in the forward process of each MeteoRA module through all LLM’s decoder blocks. MeteoRA also applies top- and top- gating strategies as detailed in Appendix A.1.
@@ 4 Evaluation
We conduct experiments on individual and composite tasks as detailed in Section 4.1. For our base models, we use two well-known LLMs, LlaMA2-13B (Touvron et al., 2023) and LlaMA3-8B (Meta, 2024). The code and the models are available222The implementation code is accessible at https://github.com/ParagonLight/meteor-of-lora, and the two MeteoRA embedded LLMs are available athttps://huggingface.co/ParagonLight/MeteoRA-llama2-13b and https://huggingface.co/ParagonLight/MeteoRA-llama3-8b.
@@ 4.1 Evaluation settings
Metrics. We apply a zero-shot evaluation setting for all tasks, adding brief task descriptions for tasks such as CNN/DailyMail and the three translation tasks that do not inherently include task descriptions. As for metrics, we use accuracy for multiple-choice tasks and GSM8K while employing metrics such as BLEU, ROUGE-1, ROUGE-2, and ROUGE-L for other tasks.
Models. We use LlaMA2-13B and LlaMA3-8B as the base LLMs for LoRA and MeteoRA adaption. Both LlaMA models are pretrained LLMs and do not include the process of instruction tuning. We train specific LoRA adapters for each task using their respective training sets. The process of training LoRA adapters could be offline or dismissed when off-the-shelf LoRA is accessible. Then, the Gating networks, which embed the adapters in the MeteoRA module, are fine-tuned efficiently based on the balanced dataset containing 1,000 samples for each task. The Gating networks for 28 tasks take no more than 10 hours to reach the convergence with 4 H800 training via Accelerate (Gugger et al., 2022). For scenarios where the training data for the original LoRA adapter is limited, we train Gating networks using a top-2 strategy, with only 100 and 5 samples accessible per task.
For baseline comparisons, we train one LoRA adapter (i.e., LoRA-F) using a mixed training set from all 28 tasks, and another LoRA adapter (i.e., LoRA-B) with the balanced dataset designed for training the Gating network. We also use Huggingface PEFT (short in PEFT) loading all 28 LoRA adapters (same ones used for MeteoRA) with explicit LoRA activation information during evaluation as a reference model. Additionally, we include several LoRA merge methods for comparison, including: averaging 28 LoRA adapters (referred to as Avg Merge), TIES (Yadav et al., 2024a), DARE (Yu et al., 2024), Arrow (Ostapenko et al., 2024), and LoraHub (Huang et al., 2023).
All LoRA adapters interact with all seven linear layers in LLaMA’s Decoder layer, configured with , , and a learning rate of . Due to some tasks having small training sets, the batch size for fine-tuning is set to . All our experiments were conducted on a GPU server with five H800 80G GPUs. Notice that we carefully selected the training hyperparameters for the LoRA-F and LoRA-B to ensure that their performance on the 28 tasks would not be excessively incomparable.
@@ 4.2 Main results
Figures 3(a) and 3(b) demonstrate the performance of the MeteoRA models, LoRA-F, LoRA-B, 5 LoRA merge methods, and a reference model PEFT based on LlaMA2-13B and LlaMA3-8B, respectively, across the selected 28 tasks. Table 1 shows the averaged scores in various matrics for all methods.
@@
LoRA-F | 0.730 / 0.767 | 41.27 / 42.93 | 0.318 / 0.327 | 0.136 / 0.157 | 0.294 / 0.306 |
LoRA-B | 0.666 / 0.750 | 37.98 / 38.47 | 0.314 / 0.343 | 0.128 / 0.171 | 0.288 / 0.321 |
Avg Merge | 0.370 / 0.427 | 19.23 / 39.89 | 0.231 / 0.200 | 0.082 / 0.060 | 0.184 / 0.158 |
TIES | 0.388 / 0.441 | 47.28 / 34.66 | 0.195 / 0.199 | 0.055 / 0.059 | 0.151 / 0.158 |
DARE | 0.332 / 0.404 | 46.531 / 36.74 | 0.192 / 0.188 | 0.054 / 0.056 | 0.144 / 0.147 |
Arrow | 0.569 / 0.647 | 41.03 / 29.93 | 0.281 / 0.283 | 0.123 / 0.142 | 0.234 / 0.242 |
LoraHub | 0.307 / 0.235 | 13.43 / 10.11 | 0.158 / 0.141 | 0.049 / 0.035 | 0.124 / 0.104 |
MeteoRA (T1-1k) | 0.755 / 0.811 | 36.73 / 45.64 | 0.336 / 0.338 | 0.160 / 0.158 | 0.313 / 0.314 |
MeteoRA (T2-1k) | 0.748 / 0.806 | 38.97 / 44.98 | 0.336 / 0.337 | 0.161 / 0.158 | 0.314 / 0.313 |
MeteoRA (T2-100) | 0.758 / 0.783 | 39.44 / 39.90 | 0.331 / 0.309 | 0.159 / 0.139 | 0.281 / 0.256 |
MeteoRA (T2-5) | 0.740 / 0.773 | 38.37 / 40.12 | 0.328 / 0.299 | 0.156 / 0.131 | 0.277 / 0.246 |
The evaluation results indicate that, regardless of the base LLM, the MeteoRA models utilizing the top-1 strategy achieve performance very close to the reference model PEFT, while no explicit LoRA activation/deactivation is required in MeteoRA. Although LLMs with both LoRA-F and LoRA-B reach comparable performance on several certain tasks, they exhibit significantly poorer outcomes on others. Additionally, MeteoRA employing the top-2 strategy, despite occasionally showing greater capability loss compared to MeteoRA with top-1 strategy, occasionally outperforms PEFT with adapters trained directly on the individual tasks. This suggests that the component in the loss function (Equation 4) becomes influential in these cases, indicating a beneficial mix of LoRA adapters from various tasks for future study. For the MeteoRA (T2-100) and MeteoRA (T2-5), although their performance shows a gap compared to MeteoRA 1k, they still outperform the baseline models on most metrics. This demonstrates that the Gating network can still learn to effectively utilize existing LoRA adapters with only a few examples.
@@ 4.4 Efficiency
To assess the efficiency of our novel forward pass designs using custom GPU kernel operators, we truncate samples from each test dataset of all 28 tasks. We evaluate these designs alongside four variants with the same hyperparameters: the upper-bound single-lora, the baseline loop-original, and two novel forward acceleration strategies based on : bmm-torch and bmm-triton, implemented by PyTorch and Triton respectively. Figure 5 displays the histogram of the overall root-of-runtime metric for each task and design. Additional evaluation is detailed in Appendix A.7.
@@ 5 Related Work
Multi-task fusion. Our proposed method falls into the field of LoRA adapter composition for multi-task fusion. The first category focuses on fusing the entire models. Researchers mainly study model ensembling and multi-task learning to achieve this goal. Existing works integrate the models under the setting of shared model architecture (Matena and Raffel, 2022; Jin et al., 2022; Wu et al., 2023c; Yadav et al., 2024b). Others focus on merging models with various architectures or from different tasks. Both methods (Stoica et al., 2023; Liu et al., 2022) try to merge models that are trained for various tasks without additional training. The second category is more concerned with fusion in terms of the tasks. Ilharco et al. (2022) proposes a model editing method via task vectors. Sun et al. (2022) leverage in-context learning with few-shots to enhance the performance of unseen tasks. However, these methods require multi-task training or prior knowledge for the evaluation tasks. Our method embeds off-the-shelf LoRA adapters with a Gate network in the MeteoRA module. None of the examples (zero-shot) are required for all individual tasks.
@@ A.4 Experimental results of 28 tasks
Table 4, Table 5, Table LABEL:tab:28_tasks_result_llama2_non_choice and Table LABEL:tab:28_tasks_result_llama3_non_choice show the detailed evaluation results of different models on the 28 selected tasks. When drawing Figure 3, for tasks we use BLEU and ROUGE as metrics, we selected BLEU for news_commentary_de, news_commentary_es, and news_commentary_it, while opting for ROUGE-L for the remaining tasks.
@@
Task Name | PEFT (reference) | LoRA-F | LoRA-B | Avg LoRA | TIES | DARE | Arrow | LoraHub | MeteoRA (T1-1k) | MeteoRA (T2-1k) | MeteoRA(T2-100) | MeteoRA(T2-5) |
AbsNarr | 0.863 | 0.758 | 0.720 | 0.562 | 0.340 | 0.190 | 0.788 | 0.278 | 0.858 | 0.860 | 0.860 | 0.868 |
ConParaKC | 0.999 | 0.999 | 0.994 | 0.424 | 0.579 | 0.554 | 0.836 | 0.514 | 0.999 | 0.999 | 0.999 | 0.998 |
CSAlg | 0.841 | 0.848 | 0.818 | 0.333 | 0.504 | 0.572 | 0.712 | 0.515 | 0.841 | 0.818 | 0.826 | 0.826 |
DisflQA | 0.690 | 0.670 | 0.573 | 0.306 | 0.356 | 0.307 | 0.506 | 0.236 | 0.679 | 0.684 | 0.683 | 0.661 |
ElemMath | 0.801 | 0.671 | 0.375 | 0.707 | 0.249 | 0.212 | 0.369 | 0.364 | 0.794 | 0.725 | 0.771 | 0.718 |
EpiReason | 1.000 | 1.000 | 0.995 | 0.367 | 0.390 | 0.367 | 0.685 | 0.233 | 1.000 | 0.998 | 1.000 | 1.000 |
FormFall | 0.999 | 0.921 | 0.565 | 0.510 | 0.510 | 0.510 | 0.961 | 0.299 | 0.999 | 0.996 | 1.000 | 0.999 |
GSWiki | 0.906 | 0.877 | 0.842 | 0.639 | 0.646 | 0.591 | 0.839 | 0.260 | 0.887 | 0.872 | 0.879 | 0.881 |
GSM8K | 0.458 | 0.428 | 0.338 | 0.062 | 0.058 | 0.052 | 0.252 | 0.155 | 0.420 | 0.439 | 0.427 | 0.397 |
LangID | 0.874 | 0.728 | 0.542 | 0.235 | 0.403 | 0.283 | 0.455 | 0.253 | 0.872 | 0.854 | 0.869 | 0.848 |
LogDeduc | 0.720 | 0.653 | 0.680 | 0.330 | 0.360 | 0.323 | 0.587 | 0.473 | 0.713 | 0.717 | 0.720 | 0.723 |
ObjCount | 0.740 | 0.690 | 0.725 | 0.330 | 0.285 | 0.245 | 0.290 | 0.180 | 0.735 | 0.725 | 0.740 | 0.720 |
ParaSeg | 0.214 | 0.274 | 0.214 | 0.047 | 0.050 | 0.036 | 0.178 | 0.015 | 0.195 | 0.182 | 0.297 | 0.295 |
PlayDiag | 0.649 | 0.649 | 0.650 | 0.649 | 0.649 | 0.649 | 0.649 | 0.265 | 0.649 | 0.649 | 0.649 | 0.649 |
QuestSel | 0.927 | 0.801 | 0.794 | 0.509 | 0.617 | 0.506 | 0.937 | 0.291 | 0.937 | 0.934 | 0.924 | 0.934 |
ColorReason | 0.950 | 0.950 | 0.950 | 0.400 | 0.400 | 0.393 | 0.660 | 0.515 | 0.930 | 0.940 | 0.935 | 0.810 |
StratQA | 0.731 | 0.729 | 0.722 | 0.367 | 0.606 | 0.558 | 0.707 | 0.573 | 0.742 | 0.722 | 0.718 | 0.722 |
TrackObj | 0.188 | 0.181 | 0.188 | 0.191 | 0.101 | 0.103 | 0.181 | 0.125 | 0.173 | 0.192 | 0.185 | 0.195 |
UnitConv | 0.755 | 0.779 | 0.707 | 0.358 | 0.370 | 0.274 | 0.534 | 0.308 | 0.727 | 0.735 | 0.729 | 0.604 |
VitaFact | 0.899 | 0.908 | 0.812 | 0.171 | 0.640 | 0.200 | 0.817 | 0.245 | 0.897 | 0.897 | 0.897 | 0.893 |
WinoWhy | 0.802 | 0.797 | 0.767 | 0.002 | 0.028 | 0.038 | 0.005 | 0.344 | 0.797 | 0.767 | 0.801 | 0.795 |
Average | 0.762 | 0.729 | 0.665 | 0.357 | 0.388 | 0.332 | 0.569 | 0.307 | 0.754 | 0.748 | 0.758 | 0.740 |